OpenStudy (anonymous):

((1/x+6)^2)+(4/x+6)=5

OpenStudy (anonymous):

Never mind figured it out

jhonyy9 (jhonyy9):

if (x+6) is denominator so than will get 1+4(x+6)=5(x+6)2 where (x+6)2 is x+6 squared 1+4x+24=5(x2+12x+36) 4x+25=5x2+60x+180 5x2+56x+155=0

jhonyy9 (jhonyy9):

now you need solve this quadratic equation